# Interaction of the Primitive Equations with Sea Ice Dynamics

### v3

@@ 7.2. The hydrostatic Dirichlet-to-Neumann operator
It follows by Subsection 5.1 that $-\mathrm{A}^{\mathrm{H}}+\omega$ has bounded imaginary powers, see also (8.7). We deduce therefrom that $D((-\mathrm{A}^{\mathrm{H}})^{\delta})\hookrightarrow\mathrm{H}_{\mathrm{per}}^{{2\delta},q}(G)^{2}$ and then conclude the following result.
[…]
@@ 8.1. Maximal regularity and bounded $\mathcal{H}^{\infty}$-calculus of the linearized operator matrix
Fix $(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$ and consider the operator matrix $\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$. Throughout this section, we omit the fixed variable $(v_{\mathrm{ice},0},h_{0},a_{0})$ and denote the operator matrix by $\mathcal{A}:=\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$. However, we impose the required regularity conditions on $(v_{\mathrm{ice},0},h_{0},a_{0})$ when necessary. Note that $\mathcal{A}$ does not have a diagonal domain. To deal with this issue, we follow the ideas from [21, 6, 8]. We introduce the operator matrix $\tilde{\mathcal{A}}\colon D(\tilde{\mathcal{A}})\subset\mathcal{X}_{0}\to\mathcal{X}_{0}$ given by
$\tilde{\mathcal{A}}$ $:=\begin{pmatrix}\mathrm{A}^{\mathrm{atm}}&0&0&0&0\\ 0&\mathrm{A}^{\mathrm{ocn}}_{0}+\mathrm{L}_{0}C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}&-\mathrm{L}_{0}(\mathrm{A}^{\mathrm{H}}-C_{\mathrm{o,i}}\mathrm{N}_{0})&0&0\\ 0&-C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}&\mathrm{A}^{\mathrm{H}}-C_{\mathrm{o,i}}\mathrm{N}_{0}&\mathrm{B}_{h}&\mathrm{B}_{a}\\ 0&0&0&d_{\mathrm{h}}\Delta_{\mathrm{H}}&0\\ 0&0&0&0&d_{\mathrm{a}}\Delta_{\mathrm{H}}\end{pmatrix},$
$D(\tilde{\mathcal{A}})$ $:=D(\mathrm{A}^{\mathrm{atm}})\times D(\mathrm{A}^{\mathrm{ocn}}_{0})\times D(\mathrm{A}^{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\times D(\Delta_{\mathrm{H}}).$
@@
(8.1) $\mathcal{X}_{1}$ =$:\displaystyle=\{v\in D(A^{\mathrm{atm}})\times D(A^{\mathrm{ocn}})\times D(\mathrm{A}^{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\times D(\Delta_{\mathrm{H}})$ :
@@
[…]
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\frac{1}{2}$ and $(v_{\mathrm{atm},0},v_{\mathrm{ocn},0},v_{0})\in V$. Then there exists a constant $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, $-\mathrm{J}_{1}+\omega$ admits bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}$.
@@
with diagonal domain. We now verify the assumptions of Subsection 8.1 for $\tilde{\mathrm{J}}_{1}$. The diagonal dominance of $\tilde{\mathrm{J}}_{1}$ follows from Subsection 7.1, as the latter result also implies that $\mathrm{C}\in\mathcal{L}((D(-\mathrm{A})^{\gamma}),\mathrm{L}^{q}(G)^{2})$ for all $\gamma\in(\nicefrac{{1}}{{2}}(1+\nicefrac{{1}}{{q}}),1]$. It remains to verify (8.5). Let $\delta\in(0,\nicefrac{{1}}{{2q}})$. As the bounded $\mathcal{H}^{\infty}$-calculus of $-\mathrm{A}^{\mathrm{H}}+\omega$ and $-\mathrm{A}^{\mathrm{ocn}}_{0}$ from Subsection 5.1 and Subsection 6.1 especially yields the boundedness of the imaginary powers of these operators, see also the relation (8.7) below, it follows that
@@
Now, Subsection 8.1 implies that $-\tilde{\mathrm{J}}_{1}+\omega$, given by (8.6), admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}$ for some constant $\omega\in\mathbb{R}$. By Subsection 7.1 and Subsection 7.1, the operator $\mathrm{L}_{0}C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}$ is relatively $(-\mathrm{A}^{\mathrm{ocn}})^{\gamma}$-bounded for all $\gamma\in(\nicefrac{{1}}{{2}}(1+\nicefrac{{1}}{{q}}),1]$. In addition, by Subsection 7.1 and Subsection 7.2, the operators $\mathrm{L}_{0}C_{\mathrm{o,i}}\mathrm{N}_{0}$ and $C_{\mathrm{o,i}}\mathrm{N}_{0}$ are relatively $(-\mathrm{A}^{\mathrm{H}})^{\theta}$-bounded for all $\theta\in(\nicefrac{{1}}{{2}},1)$. Therefore, the claim follows from perturbation theory for the $\mathcal{H}^{\infty}$-calculus, see for instance [30, Proposition 13.1]. ∎
@@
[…]
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ and $(v_{\mathrm{atm},0},v_{\mathrm{ocn},0},v_{0})\in V$. Then there exists a constant $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, the operator matrix $-\mathcal{A}+\omega$, with $\mathcal{A}$ given in (4.4), admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$.
@@
In the sequel, the constants $\omega\in\mathbb{R}$ may change from line to line. It follows from [39, Theorem 8.22] that $-\mathrm{J}_{2}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$. Combining this with Subsection 8.1, we deduce that $-\mathrm{diag}({\mathrm{J}_{1},\mathrm{J}_{2}})+\omega$ with diagonal domain admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}\times\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$. Further, we know from Subsection 5.1 that $\mathrm{B}^{\prime}$ is relatively $(-\mathrm{A}^{\mathrm{H}}+\tilde{\omega})^{\frac{1}{2}}$-bounded, and it follows from [30, Proposition 13.1] and (8.3) that $-\mathrm{J}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}\times\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$.
Invoking Subsection 6.1(a), we conclude from (8.2) that $-\tilde{\mathcal{A}}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$. The claim finally follows from Subsection 8.1 together with the fact that the bounded $\mathcal{H}^{\infty}$-calculus is preserved under similarity transforms, see for instance [13, Proposition 2.11]. ∎
@@
    The operator $-\mathcal{A}+\omega$ has an $\mathcal{R}$-bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$ with angle $\Phi_{-\mathcal{A}+\omega}<\nicefrac{{\pi}}{{2}}$.
@@
    It holds that $-\mathcal{A}+\omega\in\mathcal{BIP}(\mathcal{X}_{0})$, i. e., the operator $-\mathcal{A}+\omega$ has bounded imaginary powers.
@@
    The operator $-\mathcal{A}+\omega$ has the property of maximal $\mathrm{L}^{p}[0,\infty)$-regularity in $\mathcal{X}_{0}$.
@@
For (a), we observe that the space $\mathcal{X}_{0}$ has the property $(\alpha)$ as a Cartesian product of closed subspaces of $\mathrm{L}^{q}$-spaces. By [28, Theorem 5.3], $\mathcal{R}$-boundedness of the $\mathcal{H}^{\infty}$-calculus is then equivalent to the $\mathcal{H}^{\infty}$-calculus and the angles coincide, showing (a).
@@
see for example [13, Section 4.4], and the observation that $\mathcal{X}_{0}$ is in particular a UMD-space.
The first part of property (d) follows from [13, Theorem 2.5]. Note that for $\tilde{\mathcal{A}}$, we obtain from the bounded imaginary powers of Hibler’s operator and the hydrostatic Stokes operators in view of Subsection 5.1 and Subsection 6.1 as well as (8.7) that the complex interpolation spaces are given by
$\tilde{\mathcal{X}_{\beta}}$ $:=[\mathcal{X}_{0},D(-\tilde{\mathcal{A}})]_{\beta}$
$=(\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{atm}})\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}))\times(\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{ocn}})\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}))\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)^{2}\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G),$
where for the interpolation spaces with periodic boundary conditions, we refer to [25, Section 4]. Now, the isomorphism $\mathcal{S}$ from Subsection 8.1 is also an isomorphism in the category of Banach couples, see [2, Section I.2.1]. Thanks to the functoriality of the interpolation, the complex interpolation spaces for $\mathcal{A}$ are given by
@@ 9. Local Well-Posedness
It follows from Subsection 8.1(c) that $-\mathcal{A}_{\omega}$ has the property of maximal $\mathrm{L}^{p}$-regularity on $\mathcal{X}_{0}$, and Subsection 8.2 yields that suitable Lipschitz estimates are valid for $\mathcal{A}_{\omega}$ and $\mathcal{F}_{\omega}$. The assertion then follows from the local existence theorem for quasilinear evolution equations, see [41, Theorem 5.1.1], and the fact that finding a solution to the quasilinear abstract Cauchy problem (4.6) is equivalent to solving the coupled system of PDEs (3.4) by the reformulation in Section 4. ∎

### v4

@@ 7.2. The hydrostatic Dirichlet-to-Neumann operator
It follows by Subsection 5.1 that $\mathrm{A}^{\mathrm{H}}+\omega$ has bounded imaginary powers, see also (8.7). We deduce therefrom that $D((-\mathrm{A}^{\mathrm{H}})^{\delta})\hookrightarrow\mathrm{H}_{\mathrm{per}}^{{2\delta},q}(G)^{2}$ and then conclude the following result.
Corollary 7.10.
@@ 8.1. Maximal regularity and bounded $\mathcal{H}^{\infty}$-calculus of the linearized operator matrix
Fix $(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$ and consider the operator matrix $\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$. Throughout this section, we omit the fixed variable $(v_{\mathrm{ice},0},h_{0},a_{0})$ and denote the operator matrix by $\mathcal{A}:=\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$. However, we impose the required regularity conditions on $(v_{\mathrm{ice},0},h_{0},a_{0})$ when necessary. Note that $\mathcal{A}$ does not have a diagonal domain. To deal with this issue, we follow the ideas from [6]. We introduce the operator matrix $\tilde{\mathcal{A}}\colon D(\tilde{\mathcal{A}})\subset\mathcal{X}_{0}\to\mathcal{X}_{0}$ given by
$\tilde{\mathcal{A}}$ $=\begin{pmatrix}\mathrm{A}^{\mathrm{atm}}&0&0&0&0\\ 0&\mathrm{A}^{\mathrm{ocn}}_{0}+\mathrm{L}_{0}C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}&-\mathrm{L}_{0}(\mathrm{A}^{\mathrm{H}}-C_{\mathrm{o,i}}\mathrm{N}_{0})&0&0\\ 0&-C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}&\mathrm{A}^{\mathrm{H}}-C_{\mathrm{o,i}}\mathrm{N}_{0}&\mathrm{B}_{h}&\mathrm{B}_{a}\\ 0&0&0&d_{\mathrm{h}}\Delta_{\mathrm{H}}&0\\ 0&0&0&0&d_{\mathrm{a}}\Delta_{\mathrm{H}}\end{pmatrix}$,
[…]
@@
(8.1) $\mathcal{X}_{1}$ $=\{v\in D(A^{\mathrm{atm}})\times D(A^{\mathrm{ocn}})\times D(\mathrm{A}^{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\times D(\Delta_{\mathrm{H}}):$
@@
Lemma 8.3.
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\frac{1}{2}$ and $(v_{\mathrm{atm},0},v_{\mathrm{ocn},0},v_{0})\in V$. Then there exists a constant $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, $\mathrm{J}_{1}+\omega$ admits bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}$.
@@
with diagonal domain. We now verify the assumptions of Subsection 8.1 for $\tilde{\mathrm{J}}_{1}$. The diagonal dominance of $\tilde{\mathrm{J}}_{1}$ follows from Subsection 7.1, as the latter result also implies that $\mathrm{C}\in\mathcal{L}((D(-\mathrm{A})^{\gamma}),\mathrm{L}^{q}(G)^{2})$ for all $\gamma\in(\nicefrac{{1}}{{2}}(1+\nicefrac{{1}}{{q}}),1]$. It remains to verify (8.5). Let $\delta\in(0,\nicefrac{{1}}{{2q}})$. As the bounded $\mathcal{H}^{\infty}$-calculus of $\mathrm{A}^{\mathrm{H}}+\omega$ and $\mathrm{A}^{\mathrm{ocn}}_{0}$ from Subsection 5.1 and Subsection 6.1 especially yields the boundedness of the imaginary powers of these operators, see also the relation (8.7) below, it follows that
@@
Now, Subsection 8.1 implies that $\tilde{\mathrm{J}}_{1}+\omega$, given by (8.6), admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}$ for some constant $\omega\in\mathbb{R}$. By Subsection 7.1 and Subsection 7.1, the operator $\mathrm{L}_{0}C_{\mathrm{o,i}}\partial_{\nu_{\Gamma_{o}}}$ is relatively $(-\mathrm{A}^{\mathrm{ocn}})^{\gamma}$-bounded for all $\gamma\in(\nicefrac{{1}}{{2}}(1+\nicefrac{{1}}{{q}}),1]$. In addition, by Subsection 7.1 and Subsection 7.2, the operators $\mathrm{L}_{0}C_{\mathrm{o,i}}\mathrm{N}_{0}$ and $C_{\mathrm{o,i}}\mathrm{N}_{0}$ are relatively $(-\mathrm{A}^{\mathrm{H}})^{\theta}$-bounded for all $\theta\in(\nicefrac{{1}}{{2}},1)$. Therefore, the claim follows from perturbation theory for the $\mathcal{H}^{\infty}$-calculus, see for instance [27, Proposition 13.1]. ∎
@@
Proposition 8.4.
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ and $(v_{\mathrm{atm},0},v_{\mathrm{ocn},0},v_{0})\in V$. Then there exists a constant $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, the operator matrix $\mathcal{A}+\omega$, with $\mathcal{A}$ given in (4.4), admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$.
@@
In the sequel, the constants $\omega\in\mathbb{R}$ may change from line to line. It follows from [34, Theorem 8.22] that $\mathrm{J}_{2}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$. Combining this with Subsection 8.1, we deduce that $\mathrm{diag}({\mathrm{J}_{1},\mathrm{J}_{2}})+\omega$ with diagonal domain admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}\times\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$. Further, we know from Subsection 5.1 that $\mathrm{B}^{\prime}$ is relatively $(-\mathrm{A}^{\mathrm{H}}+\tilde{\omega})^{\frac{1}{2}}$-bounded, and it follows from [27, Proposition 13.1] and (8.3) that $\mathrm{J}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{2}\times\mathrm{L}^{q}(G)\times\mathrm{L}^{q}(G)$ for some $\omega\in\mathbb{R}$.
Invoking Subsection 6.1(a), we conclude from (8.2) that $\tilde{\mathcal{A}}+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$. The claim finally follows from Subsection 8.1 together with the fact that the bounded $\mathcal{H}^{\infty}$-calculus is preserved under similarity transforms, see for instance [15, Proposition 2.11]. ∎
@@
    The operator $\mathcal{A}+\omega$ has an $\mathcal{R}$-bounded $\mathcal{H}^{\infty}$-calculus on $\mathcal{X}_{0}$ with angle $\Phi_{-\mathcal{A}+\omega}<\nicefrac{{\pi}}{{2}}$.
@@
    It holds that $\mathcal{A}+\omega\in\mathcal{BIP}(\mathcal{X}_{0})$, i. e., the operator $\mathcal{A}+\omega$ has bounded imaginary powers.
@@
    The operator $\mathcal{A}+\omega$ has the property of maximal $\mathrm{L}^{p}[0,\infty)$-regularity in $\mathcal{X}_{0}$.
@@
For (a), we observe that the space $\mathcal{X}_{0}$ has the property $(\alpha)$. Thus $\mathcal{R}$-boundedness of the $\mathcal{H}^{\infty}$-calculus is then equivalent to the $\mathcal{H}^{\infty}$-calculus and the angles coincide, showing (a).
@@
see for example [15, Section 4.4], and the observation that $\mathcal{X}_{0}$ is in particular a UMD-space.
The first part of property (d) follows from [15, Theorem 2.5]. Note that for $\tilde{\mathcal{A}}$, we obtain from the bounded imaginary powers of Hibler’s operator and the hydrostatic Stokes operators in view of Subsection 5.1 and Subsection 6.1 as well as (8.7) that the complex interpolation spaces are given by
$\tilde{\mathcal{X}_{\beta}}$ $=[\mathcal{X}_{0},D(-\tilde{\mathcal{A}})]_{\beta}$
$=(\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{atm}})\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}))\times(\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{ocn}})\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}))\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)^{2}\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G).$
Now, the isomorphism $\mathcal{S}$ from Subsection 8.1 is also an isomorphism in the category of Banach couples, see [2, Section I.2.1]. Thanks to the functoriality of the interpolation, the complex interpolation spaces for $\mathcal{A}$ are given by
@@ 9. Local Well-Posedness
It follows from Subsection 8.1(c) that $\mathcal{A}_{\omega}$ has the property of maximal $\mathrm{L}^{p}$-regularity on $\mathcal{X}_{0}$, and Subsection 8.2 yields that suitable Lipschitz estimates are valid for $\mathcal{A}_{\omega}$ and $\mathcal{F}_{\omega}$. The assertion then follows from the local existence theorem for quasilinear evolution equations, see [35, Theorem 5.1.1], and the fact that finding a solution to the quasilinear abstract Cauchy problem (4.6) is equivalent to solving the coupled system of PDEs (3.4) by the reformulation in Section 4. ∎